SavelyevA4523:

Sophie works 4 hours a day, and earns a salary of $500.00 per month. She earns $10 for every extra hour that she works. If she earned $650.00 last month, how many extra hours did Sophie work?

SavelyevA4523:

15

Hero:

x represents extra hours worked therefore 500 + 10x = 650 10x = 650 - 500 10x=150 x=150/10 x =15 (extra hours) you are correct.

Hero:

@SavelyevA4523

SavelyevA4523:

Thank You very much @Hero

Hero:

You're welcome